# On long-duration storage, weather uncertainty and limited foresight

### v1

@@ 3.3 Scenarios
The comparison between Limited Foresight and Perfect Foresight is conducted for different countries, namely Germany, Spain and United Kingdom (UK), and for different assumptions with respect to the availability of a backup technology/backup supply.
First, we choose three different countries to reflect different empirical distributions of wind and solar PV capacity factors as well as heating demand.
@@
Scenario Cost of outside option Maximum
@@ 3.4 Implementation
While SDDP has been shown to converge almost surely in finitely many iterations [philpott_convergence_2008], practical termination criteria remain an open research topic [dowson_policy_2020], as the upper bound is probabilistic. We opt for the aforementioned time and iteration limits and inspect convergence ex-post. In the Section SI.7, we show that capacity decisions begin to stabilize after around 3000 iterations, see Figure SI.1. Nevertheless, subsequent iterations serve to refine the policy, which is particularly useful in order to obtain granular LDES bidding curves (dual policy). We report additional details on convergence for each model in Table SI.1. We facilitate a comparison to the Perfect Foresight model by simulating the 35 historical weather years $\{1982/83,\dots,2016/17\}$ in our sample using the trained policy $\{\hat{\mathcal{Q}}\}$. That is, we solve the dispatch for a given weather year by solving each monthly subproblem with the learned expected cost-to-go functions.
@@ 4.1 LDES stockpiles against bad system states
Fig. 4 shows the the distribution (5th-95th, 25th-75th percentile and mean trajectories as well as the actual paths) of storage trajectories when simulating the Perfect Foresight case (blue; PF) and the Limited Foresight case (red; LF) over the 35 historical weather years between July 1982 and June 2017. These results are for the No Imports scenario. Fig. SI.3 and Fig. SI.4 present the trajectories for the other two scenarios.
Figure 4: LDES storage level trajectories: Panels show distribution of LDES storage level trajectories over 35 historical years (July 1982-Jun 2017) for Germany (top), Spain (middle) and United Kingdom (bottom). Red components refer to Limited Foresight and blue components refer to Perfect foresight. Trajectories are normalized by the endogenous storage target level. The thick lines show the mean trajectory. Darker shaded areas represent the 25th-75th percentiles, lighter areas show the 5th-95th percentile of the distribution. Thin lines show the actual trajectories.
@@
For all three countries the mean trajectory in the LF case lies well above that of PF. This difference is largest in December when LF LDES levels are still close to full capacity whereas PF LDES levels cross the zero line entering into a deficit vis a vis the target level. The LF model reaches mean levels close to full capacity as early as October and maintains them until the turn of the year in the case of United Kingdom. This way, the LDES operator hedges against bad weather states with prolonged periods of VRE scarcity potentially coinciding with cold spells ([schmidt_european_2025]). In the No Imports scenario such states could become very costly as load would have to be shed at the cost of VOLL if the LDES cannot supply during a positive net load event. By contrast, the PF model can foresee all periods of scarcity and abundance over the model horizon and behave much more aggressively. In PF the maximum dispatch event across all weather years defines the storage size [kittel_coping_2025] and only then full capacity is met. As Fig. SI.3 and Fig. SI.4 show, the difference between the trajectory distribution dissipates as other backup options in the form of H2 spot imports become available.
@@
Apart from the common dynamics across countries, there are also stark differences. The Spanish trajectories under LF are fairly close to those under PF. The LF mean trajectory enters the deficit roughly one month after its PF counterpart but runs a deficit of up 6.5 TWh before returning to the target level. For Spain, both distributions, PF and LF, are rather narrow. The United Kingdom presents a different picture. The LF mean trajectory never enters into a deficit. The maximum deficit across all years is just 3.9 TWh compared to 20 TWh under PF. Both distributions are rather wide. The PF distribution begins running wide much earlier than the LF distribution.
@@
We present an alternative perspective in Figure SI.5, which gives the storage trajectories in their chronological order over the entire period from 1982 to 2017 and at absolute scale. Spain’s trajectories are far more even and the difference between LF and PF is rather stable over the sample with the LF LDES consistently reaching its low around 10 TWh and the PF LDES frequently almost emptying the storage. By contrast, the United Kingdom’s trajectory is far more erratic. While the storage level fluctuates wildly over the years in the PF case, it consistently charges to maximum capacity in preparation for the winter in LF. Interestingly, the PF storage level exceeds the LF storage level in few instances for Germany and the UK enabling deeper dispatch episodes, especially around the Dunkelflaute of 1996/97 and the cold winter of 2010/11 in Britain.
@@
Spain and United Kingdom represent solar-dominated and wind-dominated systems respectively. As evident from Fig. SI.14 the empirical distribution of wind capacity factors is wide and wind tends to follow synoptic patterns [schlott_impact_2018]. By contrast, seasonal and diurnal patterns for solar generation are much more stable and the distribution tends to be narrow by comparison. The width of the PF distribution for the UK reflects sporadic dispatch in windless periods and charging in windy periods also occurring in the winter period. Lacking perfect foresight, the LF model cannot rely on windy winter periods and stockpiles in expectation of extreme winter periods. In Spain, the seasonality is much more pronounced reflecting predictably lower solar capacity factors in winter. Germany, neither wind- nor solar-dominated to the same degree, exhibits both patterns.
@@
As the costs of bad system states reduce as spot imports are allowed in scenarios Constrained Imports and Unlimited Imports, the differences between the two models dissipate as the optimal policy replaces some stockpiling by spot imports, see Figure SI.3 and Figure SI.4.
@@ 4.2 System value of solar PV increases under uncertainty
The observed differences in LDES behavior also have an effect on optimal capacity decisions. Again we focus on the results in the No Imports scenario. As Figure 5 shows the differences between PF (blue squares) and LF (red circles) are modest compared to the wide range observed when optimizing capacities for each weather year separately (type A from Section 2), as indicated by the gray crosses. The latter is especially wide for cavern storage capacity reflecting LDES’ sensitivity to interannual weather variability.
In terms of generation capacities, we observe a significant increase in solar PV capacity between PF and LF. For Germany the increase is 48 GW or 12.8%. For the UK the increase is ca. 30 GW or 25.4%. By contrast, Spain’s PV capacity only increases by 7.4% or 15 GW. Wind onshore capacity differences are smaller. Only for the UK we observe a larger reduction of 7.9 GW or 20%. Wind offshore capacities remain at the upper limit for Germany defined in Section 3.2. For UK, they increase slightly (9%). Having limited offshore potential, Spain’s capacities increase from 12 to 17 GW. Run-of-river, nuclear and biomass are either fixed by design or unchanged between the scenarios.
@@
The impact of limited foresight on LDES capacities is rather small. PEM electrolysis capacities remain at par with PF in Germany and UK. In Spain they increase slightly by 3 GW. In Germany and UK cavern storage capacities decrease (DE: -8%, UK: - 11%), while they increase for Spain, albeit on a lower level.
@@
The described capacity differences between PF and LF are consistent with the differences in LDES behavior discussed in the previous section. Stockpiling requires reliable electrolysis activity during the summer and autumn months, including the period leading up to the storage target, to ensure high storage levels for bad system states that are more likely to occur in winter. Solar capacity factors have a much lower variance than wind capacity factors (Figure SI.14) and therefore offer more reliable feed-in in the lead-up of the winter.
@@
While under perfect foresight, the system operator can capitalize on particularly windy periods during winter that make up for periods of extreme energy scarcity, the LF model cannot leverage these episodes to replenish LDES levels as they are highly uncertain. Accordingly, solar generation replaces some wind onshore generation (Figure SI.8), electrolysis activity more closely resembles the diurnal patterns of solar PV feed-in (Figure SI.9 and Table SI.2) and is much more even across weather years, as opposed to PF, where it follows year-specific wind patterns (Figure SI.10).
@@
The change in LDES capacities can be explained by the chronological storage trajectories in Figure SI.5. For Germany and the UK, PF can leverage infrequent periods of abundance to charge the storage prior to a large discharge due to a Dunkelflaute or cold spell. With perfect foresight, the benefit of this extra capacity outweighs the cost. The same does not hold under LF where many more weather realizations are possible. In Spain another effect in which the LF model keeps some insurance for extreme states that do not materialize over the historical weather year sequence dominates and the cavern size increases under LF.
@@
The change in battery capacity likely serves to the improve the full-load hours of electrolysis capacity by smoothing solar feed-in. Additionally, previous work has shown that under PF LDES may also engage in shorter-term balancing in order to reduce battery capacity needs [kittel_coping_2025]. Under LF, the need for stockpiling may outweigh the benefit of reducing battery capacity by running existing LDES capacity for short-term balancing.
@@
Unsurprisingly, the described effects tend to be largest for a wind-dominated system like the UK and smallest for a solar-dominated system like Spain. Additional backup from H2 spot imports in Constrained Imports and Unlimited Imports, the differences between PF and LF converge in line with the dissipation of the stockpiling behavior discussed above.
@@
Figure 5: Comparison of capacity choices: Assuming no H2 spot imports; Left panels show generation capacities; middle panels show Li-Ion capacities; right panels show capacities for LDES components. In addition to Limited Foresight (red) and Perfect Foresight (blue), we show deterministic single-year results for comparison.
@@ 4.3 LDES bidding functions
For most possible weather realizations, the 80% storage level, and everything that will be optimally added in a given realization in the following months, will suffice to cover all positive net load events over the model horizon. In these realizations the MSV will be low. In few extreme cases, the storage level will not suffice to cover all scarcity periods. In these instances the MSV is set by the VOLL or by H2 spot import prices. The LDES bid in August at 80% reflects the average over these possible outcomes and although the probability of realizations with load shedding may be low, the assumed VOLL is so high that it drives up the average MSV. As we consider lower storage levels in the same month, the probability of load shedding or H2 spot imports increases and the associated MSV increases too.
Likewise, the probability of incurring a future bad state changes by month. Prior to the winter period the probability of a future energy shortfall at a given storage level is higher than after the worst of winter is over. Consequently, the autumn curves (yellow, orange, red and pink) have a concave shape in which high bids prevail into very high storage levels. Conversely, spring curves (blue, purple, green) are convex with low bids at relatively high storage levels.
@@
Figure 6: LDES bidding curves: Monthly LDES marginal values of storage derived from trained SDDP model differentiated by alternative H2 import options.
In summary, it is the extreme weather realizations in which the LDES cannot cover all load and the associated costs to deal with that, as well as the probability of these realizations that determine the storage bid. The probability changes with the time of year and the storage level. Naturally, these probabilities depend on the joint distribution of weather variables and the optimal dispatch given a weather realization, which are unknown (ex-ante). Using SDDP, the LF model learns extreme state probabilities reflected in the bidding functions. The bidding functions become coarser (larger step sizes) towards low storage levels, as the model is unlikely to visit this state often in the forward pass simulations.
@@ 4.4 LDES Bids under uncertainty stabilize energy-only market prices
The explicit representation of weather uncertainty and the resulting LDES bidding behavior discussed in the previous section yield a wide range of bids at which the LDES may set the price, either charging (electrolysis) or discharging (H2 turbine) effectively smoothing the duration curve. By contrast, the possibility of a bad state is not priced into the perfect foresight LDES bids reflected in electricity prices when LDES is at the margin. That is, the PF LDES bids do not include the value of storage against situations that never materialize resulting in a low MSV variation and electricity price dichotomy in the (close to) zero marginal cost systems. Figures SI.11, SI.12 and SI.13 compare the realized marginal storage values of of LF and PF.
@@ 5.1 Summary
In this work, we use a method developed for the optimization of hydro-dominated electricity systems to optimize a simplified capacity expansion model for fully renewable sector-coupled European energy systems based on wind and solar. We isolate the effect of limited foresight under weather uncertainty by comparing the output to a perfect foresight benchmark. Under weather uncertainty, LDES bidding is defined by the probability and cost of extreme system states requiring load shedding or some type of backup. The probability is conditional on the current storage level and the time of year. LDES bids therefore depend on the underlying joint weather distribution (wind, solar, hydro and heat demand).
@@
We observe indirect effects on capacity decisions. While there are only smaller changes in the LDES capacities (electrolysis, cavern storage, hydrogen gas turbine), solar PV gains in system value for providing more reliable generation than wind, especially for the production of electrolytic hydrogen in the stockpiling process. We find the highest effect in the United Kingdom where solar PV capacity increases by 25.4% and onshore wind capacity decreases by 20% under limited foresight.
Lastly, we demonstrate that LDES bidding under uncertainty critically affects price formation in fully renewable energy-only electricity markets by smoothing price duration curves as price-setting LDES bids constantly change to reflect changed probability of extreme system states.

### v2

@@ 3.3 Scenarios
The comparison between Limited Foresight and Perfect Foresight is conducted for different countries and for different assumptions with respect to the availability of backup technologies and supply.
@@
Scenario Cost of outside option Maximum quantity
@@ 3.4 Implementation
While SDDP has been shown to converge almost surely in finitely many iterations [philpott_convergence_2008], practical termination criteria remain an open research topic [dowson_policy_2020], as the upper bound is probabilistic. We opt for the aforementioned time and iteration limits and inspect convergence ex-post. In the Section SI.7, we show that capacity decisions begin to stabilize after around 3000 iterations, see Figure SI.1. Nevertheless, subsequent iterations serve to refine the policy, which is particularly useful in order to obtain granular LDES bidding curves. We report additional details on convergence for each model in Table SI.1. We facilitate a comparison to the Perfect Foresight model by simulating the 35 historical weather years $\{1982/83,\dots,2016/17\}$ in our sample using the trained policy $\{\hat{\mathcal{Q}}\}$. That is, we solve the dispatch for a given weather year by solving each monthly subproblem with the learned expected cost-to-go functions.
@@ 4.1 LDES stockpiles against bad system states
Fig. 4 shows the distribution (5th-95th, 25th-75th percentile and mean trajectories as well as the actual paths) of storage trajectories when simulating the Perfect Foresight case (blue; PF) and the Limited Foresight case (red; LF) over the 35 historical weather years between July 1982 and June 2017. These results are for the No Imports scenario. Fig. SI.3 and Fig. SI.4 present the trajectories for the other two scenarios.
Figure 4: LDES storage level trajectories: Panels show the distribution of LDES storage level trajectories over 35 historical years (July 1982-June 2017) for Germany (top), Spain (middle) and the United Kingdom (bottom). Red components refer to Limited Foresight, and blue components refer to Perfect Foresight. Trajectories are normalized by the endogenous storage target level. The thick lines show the mean trajectory. Darker shaded areas represent the 25th-75th percentiles, lighter areas show the 5th-95th percentile of the distribution. Thin lines show the actual trajectories.
@@
For all three countries, the mean trajectory in the LF case lies well above that of PF. This difference is largest in December when LF LDES levels are still close to full capacity, whereas PF LDES levels cross the zero line, entering into a deficit vis-à-vis the target level. The LF model reaches mean levels close to full capacity as early as October and maintains them until the turn of the year in the case of the United Kingdom. This way, the LDES operator hedges against bad weather states with prolonged periods of VRE scarcity potentially coinciding with cold spells [schmidt_mix_2025]. In the No Imports scenario, such states could become very costly, as load would have to be shed at the cost of VOLL if the LDES cannot supply during a positive net load event. By contrast, the PF model can foresee all periods of scarcity and abundance over the model horizon and behaves much more aggressively. Under PF, the maximum dispatch event across all weather years defines the storage size [kittel_coping_2025] and only then full capacity is met.
@@
Apart from these common dynamics across countries, there are also significant differences. The Spanish trajectories under LF are fairly close to those under PF. The LF mean trajectory enters the deficit roughly one month after its PF counterpart but runs a negative balance of up to 6.5 TWh before returning to the target level. Both distributions, PF and LF, are rather narrow. The United Kingdom presents a different picture. The LF mean trajectory never enters a deficit. The maximum deficit across all years is just 3.9 TWh compared to 20 TWh under PF. Both distributions are rather wide. The PF distribution begins running wide much earlier than the LF distribution.
@@
We present an alternative perspective in Figure SI.5, which gives the storage trajectories in their chronological order over the entire period from 1982 to 2017 and at absolute scale. Spain’s trajectories are far more even, and the difference between LF and PF is rather stable over the sample, with the LF LDES consistently reaching its low around 10 TWh and the PF LDES frequently almost emptying the storage. By contrast, the United Kingdom’s trajectory is far more erratic. While the storage level fluctuates wildly over the years in the PF case, it consistently charges to maximum capacity in preparation for the winter under LF. Interestingly, the PF storage level exceeds the LF storage level in few instances for Germany and the UK, enabling deeper dispatch episodes, especially around the Dunkelflaute of 1996/97 and the cold winter of 2010/11 in the United Kingdom.
@@
Spain and the United Kingdom represent solar-dominated and wind-dominated systems, respectively. As evident from Fig. SI.14, the empirical distribution of wind capacity factors is wide and wind tends to follow synoptic patterns [schlott_impact_2018]. By contrast, seasonal and diurnal patterns for solar generation are much more stable, and the distribution tends to be narrow in comparison. The width of the PF distribution for the UK reflects sporadic dispatch in windless periods and charging in windy periods, which also occur in the winter period. Lacking perfect foresight, the LF model cannot rely on windy winter periods and stockpiles in expectation of extreme winter periods. In Spain, the seasonality is much more pronounced, which reflects predictably lower solar capacity factors in winter. Germany, neither wind- nor solar-dominated to the same degree, exhibits both patterns.
@@
As the costs of bad system states reduce as hydrogen spot imports become available in scenarios Constrained Imports and Unlimited Imports, the differences between the two models dissipate as the optimal policy replaces some stockpiling by spot imports, see Figure SI.3 and Figure SI.4.
@@ 4.2 System value of solar PV increases under uncertainty
The observed differences in LDES behavior affect optimal capacity decisions. Again, we focus on the results in the No Imports scenario. As Figure 5 shows, the differences between PF (blue squares) and LF (red circles) are modest compared to the wide range observed when optimizing capacities for each weather year separately (type A from Section 2), as indicated by the gray crosses. The latter is especially wide for cavern storage capacity, reflecting LDES’ sensitivity to interannual weather variability.
In terms of generation capacities, we observe a significant increase in solar PV capacity between PF and LF. For Germany, the increase is 48 GW or 12.8%. For the UK, the increase is ca. 30 GW or 25.4%. By contrast, Spain’s PV capacity only increases by 7.4% or 15 GW. Wind onshore capacity differences are smaller. Only for the UK, we observe a larger reduction of 7.9 GW or 20%. Wind offshore capacities remain at the upper limit for Germany defined in Section 3.2. For the UK, they increase slightly (9%). Having limited offshore potential, Spain’s capacities increase significantly on a relative basis, from 12 to 17 GW. Run-of-river, nuclear and biomass are either fixed by design or unchanged between the scenarios.
@@
The impact of limited foresight on LDES capacities is rather small. PEM electrolysis capacities remain at par with PF in Germany and the UK. In Spain, they increase slightly by 3 GW. In Germany and the UK, cavern storage capacities decrease (DE: -8%, UK: - 11%), while they increase for Spain, albeit on a lower level.
@@
The described capacity differences between PF and LF are consistent with the differences in LDES behavior discussed in the previous section. Stockpiling requires reliable electrolysis activity during the summer and autumn months, including the period leading up to the storage target, to ensure high storage levels for bad system states that are more likely to occur in winter. Solar capacity factors have a much lower variance than wind capacity factors (Figure SI.14) and therefore offer more reliable feed-in in the lead-up to winter.
@@
While under perfect foresight, the system operator can capitalize on particularly windy periods during winter that make up for periods of extreme energy scarcity, the LF model cannot leverage these episodes to replenish LDES levels as they are highly uncertain. Accordingly, solar generation replaces some wind onshore generation (Figure SI.8), electrolysis activity more closely reflects the diurnal patterns of solar PV feed-in (Figure SI.9 and Table SI.2) and is much more even across weather years, as opposed to PF, where it follows year-specific wind patterns (Figure SI.10).
@@
The change in LDES capacities can be explained by the chronological storage trajectories in Figure SI.5. For Germany and the UK, PF can leverage infrequent periods of abundance to charge the storage prior to a large discharge due to a Dunkelflaute or cold spell. With perfect foresight, the benefit of this extra capacity outweighs the cost. The same does not hold under LF, where many more weather realizations are possible. In Spain another effect, in which the LF model keeps some insurance for extreme states that do not materialize over the historical weather year sequence, dominates and the cavern size increases under LF.
@@
The change in battery capacity likely serves to the improve the full-load hours of electrolysis capacity by smoothing solar feed-in. Additionally, previous work has shown that under PF, LDES may also engage in shorter-term balancing in order to reduce battery capacity needs [kittel_coping_2025]. Under LF, the need for stockpiling likely outweighs the benefit of reducing battery capacity by running existing LDES capacity for short-term balancing.
@@
Unsurprisingly, the described effects tend to be largest for a wind-dominated system like the UK and smallest for a solar-dominated system like Spain. With additional backup from H2 spot imports in Constrained Imports and Unlimited Imports, the differences between PF and LF converge in line with the dissipation of the stockpiling behavior discussed above.
@@
Figure 5: Comparison of capacity choices: Assuming No Imports; Left panels show generation capacities; middle panels show Li-Ion capacities; right panels show capacities for LDES components. In addition to Limited Foresight (red) and Perfect Foresight (blue), gray crosses represent deterministic single-year results for comparison.
@@ 4.3 LDES bidding functions
For most possible weather realizations, the 80% storage level, and everything that will be optimally added in a given realization in the following months, will suffice to cover all positive net load events over the model horizon. In these realizations, the MSV will be low. In a few extreme cases, the storage level will not suffice to cover all scarcity periods. In these instances, the MSV is set by the VOLL or by H2 spot import prices. The LDES bid in August at 80% reflects the average over these possible outcomes, and although the probability of realizations with load shedding may be low, the assumed VOLL is so high that it drives up the average MSV. As we consider lower storage levels in the same month, the probability of load shedding or H2 spot imports increases and the associated MSV increases too.
Likewise, the probability of incurring a future bad state changes by month. Prior to the winter period, the probability of a future energy shortfall at a given storage level is higher than after the worst of winter is over. Consequently, the autumn curves (yellow, orange, red and pink) have a concave shape in which high bids prevail at very high storage levels. Conversely, spring curves (blue, purple, green) are convex with low bids at relatively high storage levels.
@@
Figure 6: LDES bidding curves: Monthly LDES marginal values of storage derived from trained SDDP model differentiated by alternative H2 import options for Germany.
In summary, it is the probability of extreme weather realizations in which the LDES cannot cover all the load and the associated costs to deal with them that determines the LDES bid. The probability changes with the time of year and the storage level. Naturally, it depends on the joint distribution of weather variables and the optimal dispatch given a weather realization, which are unknown (ex-ante). Using SDDP, the LF model learns extreme state probabilities reflected in the bidding functions. The bidding functions become coarser (larger step sizes) towards low storage levels, as the model is unlikely to visit this state often in the forward pass simulations.
@@ 4.4 LDES Bids under uncertainty stabilize energy-only market prices
The explicit representation of weather uncertainty and the resulting LDES bidding behavior, discussed in the previous section, yield a wide range of bids at which the LDES may set the price, either charging (electrolysis) or discharging (H2 turbine), effectively smoothing the duration curve. By contrast, the possibility of a bad state is not priced into the perfect foresight LDES bids, reflected in electricity prices when LDES is at the margin. That is, the PF LDES bids do not include the value of storage against situations that never materialize, resulting in a low MSV variation and electricity price disparity in the (close to) zero marginal cost systems. Figures SI.11, SI.12 and SI.13 compare the realized marginal storage values of LF and PF.
@@ 5.1 Summary
In this work, we use a method developed for the optimization of hydro-dominated electricity systems to optimize a simplified capacity expansion model for fully renewable sector-coupled European energy systems based on wind and solar. We isolate the effect of limited foresight under weather uncertainty by comparing the output to a perfect foresight benchmark. Under weather uncertainty, LDES bidding is defined by the probability and cost of extreme system states requiring load shedding or some type of backup. The probability is conditional on the current storage level and the time of year. LDES bids therefore depend on the underlying joint weather distribution (of wind, solar, hydro and heat demand).
@@
We observe indirect effects on capacity decisions. While there are only smaller changes in the LDES capacities (electrolysis, cavern storage, hydrogen gas turbine), solar PV gains in system value for providing more reliable generation than wind, especially for the production of electrolytic hydrogen in the stockpiling process. We find the highest effect in the United Kingdom, where solar PV capacity increases by 25% and onshore wind capacity decreases by 20% under limited foresight.
Lastly, we demonstrate that LDES bidding under uncertainty critically affects price formation in fully renewable energy-only electricity markets by smoothing price duration curves as price-setting LDES bids constantly change to reflect the changed probability of extreme system states.